OpenStudy (anonymous):

solve for x: sin(x)=2*cos(2*x)

OpenStudy (anonymous):

cos(2x) = cos^2x - sin^2x = 1-2sin^2x So, you have: sin(x) = 2(1-2sin^2x). Treat sin(x) as t. You have t = 2(1-2t^2). Solve like the quadratic equation. Find values of t. Since t = sin(x), find x.

OpenStudy (anonymous):

????